# ISAC channel simulator GBSM

### v3

@@ Abstract
Integrated sensing and communication (ISAC) has been recognized as the key technology in the vision of the sixth generation (6G) era. With the emergence of new concepts in mobile communications, the channel model is the prerequisite for system design and performance evaluation. Currently, 3GPP Release 19 is advancing the standardization of ISAC channel models. Nevertheless, a unified modeling framework has yet to be established. This paper provides a simulation diagram of ISAC channel modeling extended based on the Geometry-Based Stochastic Model (GBSM), which is compatible with existing 5G channel models and the latest progress in 3GPP standardization. We first introduce the progress of the ISAC channel model standard in general. Then a concatenation channel modeling approach considering team standardization proposals is presented which is implemented on the BUPTCMCC-6G-CMG+ channel model generator. We validated the model in cumulative probability density function (CDF) in statistical extension of angle and delay, and radar cross section (RCS). Simulation results show that the proposed model can realistically characterize the feature of concatenation and RCS within the ISAC channel. At the same time, the proposed method maintains an efficient simulation.
@@
Channel model generator, concatenation, ISAC channel, RCS, 3GPP standardization.1 Introduction
@@
6G mobile communication technology is anticipated to enhance 5G communication capabilities roundly, including high capacity, low latency, and high reliability, while achieving native intelligence, ISAC, and ubiquitous connectivity [1]. The boundary between the physical and digital realms is expected to become increasingly indistinct, offering users a ubiquitous and immersive experience [2]. Concerns of communication systems will need to expand from information transmission to information acquisition and information processing[3]. ISAC is the essential technology necessary to achieve the information paradigm transformation described above. ISAC technology can facilitate the implementation of diverse use cases, particularly in localization, tracking, imaging, and the detection of target type, speed, and direction. These capabilities will enable applications in autonomous driving, industrial automation, digital twin reconstruction of physical environments, gesture detection, and activity recognition[4, 5, 6, 7, 8]. The sensing information obtained from ISAC can facilitate the development of smart homes, factories, smart cities, and intelligent transportation systems. Meanwhile, ISAC technology introduces new performance metrics relative to traditional communication technologies, including detection probability, sensing capacity density, sensing accuracy, and resolution for speed, angle, and delay [9].
@@
The transmission medium between the transmitter and receiver establishes the performance limits of communication systems. The wireless channel plays a vital role in researching and evaluating wireless communication technologies.[10, 11, 12]. The performance evaluation of the ISAC system introduces new requirements for channel models, particularly regarding sensing modes and scenarios, sensing use cases, and channel modeling methodologies. Considering sensing modes, current channel models usually utilize Bi-static modes for communication. Sensing applications necessitate considering the case of Mono-static channel models for self-sensing. Thus, ISAC channel models should incorporate both Mono-static and Bi-static sensing modes. Secondly, concerning sensing use cases involving sensing targets, ISAC applications expand the evaluation criteria to include target detection, localization, identification, and tracking. Examples encompass intrusion detection, unmanned aerial vehicle (UAV) tracking, vehicle-to-everything (V2X) communications, and the identification of diverse target types and sizes. That requires using the scattering properties of RCS, polarization characteristics, delay, and Doppler information of electromagnetic waves as they interact with targets. Consequently, channel models should integrate the RCS characteristics of various targets. Moreover, with respect to modeling methodologies, it is imperative to account for the cascaded characteristics of channels interacting with targets while coupling with existing standard communication channel models [13, 14, 15, 16].
@@
So far, the standardization of ISAC channel models by 3GPP has held five conferences from #116 to #118-bis, wherein ISAC scenarios, types of sensing targets, sensing modes, and general channel model frameworks have been defined. The defined scenarios encompass traditional urban micro (UMi), urban macro (UMa), rural macro (RMa), and indoor hotspot (InH) scenarios, in addition to V2X, UAV, and industrial Internet of things (IIoT) scenarios. Accordingly, sensing targets are classified into five categories: humans, vehicles, UAVs, automated guided vehicles (AGVs), and obstacles as shown in Table 1. Sensing modes are categorized into Bi-static and Mono-static, which are further subdivided into six specific use cases based on whether the sensing device is a user terminal (UT) or a base station (BS): BS-UT-Bi, BS-BS-Bi, UT-UT-Bi, UT-BS-Bi, UT-UT-Mono, and BS-BS-Mono [17]. In the framework of channel modeling, the ISAC channel is defined as the combination of the target channel and the background channel. The target channel includes all channel components that interact with targets, whereas the background channel comprises all other elements that do not engage with targets.
@@
For ISAC channel simulation, a channel model that separately generates communication and sensing clusters based on the 3GPP model, and utilizes sensing echoes for perception, has been proposed [18]. Some studies introduce the concept of shared clusters to explain the phenomenon of scattering body reuse between the communication and sensing channels [5, 19, 20]. Some studies focus on deterministic ISAC channel modeling [21, 22]. However, these studies do not align with the target and background channel modeling framework or the cascading channel method in 3GPP Rel-19. Most of these approaches use single-station echoes for perception and do not support Bi-static sensing modes.
To the authors’ knowledge, no unified ISAC channel simulation framework compatible with the latest 3GPP standardization progress has been proposed. Considering that standardization is primarily followed by industry and the existing standardization framework has not been fully finalized, we present our solution based on our team’s work in ISAC standardization and our understanding of standardized channel simulation. We hope this will contribute to advancing both the academic field and the establishment of standardized models. The contributions and innovations of this paper are outlined as follows:
Different from existing ISAC channel simulation models [18, 19, 20, 21, 22], we propose a unified ISAC channel simulation framework based on the 3GPP Rel-19 ISAC channel fundamental structure, where the ISAC channel is the combination of the target and background channels. This framework maximizes the use and compatibility with existing 5G standardized channel models. It incorporates our team’s contributions to the standardization process, offering a solution that aligns with the latest standardized progress and consensus in ISAC channel modeling.
The effective ISAC channel simulation method has been validated.
We compared and analyzed the down-section simulation results for the 8 cascading options presented in the RAN1 #118-bis meeting, explaining the mechanisms behind the results. Based on it, we propose a simulation method for ISAC cascading channels.
A unified RCS simulation model is introduced, which considers the progress of the 116-118bis meetings and incorporates our team’s RCS measurement results. This model applies to various target types, accounting for factors such as large-scale, small-scale, statistical, and deterministic elements.
The paper is organized as follows. Section describes the simulation architecture in detail. Section reports on evaluation metrics for ISAC. Then, Section presents the simulation results from both concatenation and RCS. Section concludes the paper.
2 Simulation Architecture
Building upon the current advancements in 3GPP standardization efforts and the GBSM simulation framework, this paper proposes an ISAC channel simulation framework designed for future standardization, which is integrated into the BUPTCMCC-6G-CMG+ channel model generator. The key characteristics of ISAC channels—including cascaded features, RCS polarization, scattering properties, and the coupling between target and background channels are considered to establish a unified model for ISAC channel simulation. In this model, the ISAC channel is divided into two components: the target channel and the background channel. The geometric schematic of the model is shown in Figure 1.
@@ 2 Simulation Architecture
The target channel is further classified into Tx-target and target-Rx sub-channels. In the Bi-static sensing mode, the target channel is constructed by cascading the Tx-target and target-Rx sub-channels, with parameters for each sub-channel generated according to the existing 3GPP statistical model. In the Mono-static sensing mode, the Tx-target link is generated using the 3GPP model, whereas the target-Rx components are derived by utilizing channel reciprocity. Furthermore, the background channel is generated directly using the 3GPP statistical channel model in the Bi-static sensing mode. As to the Mono-static sensing mode, the background channel can be derived from the 3GPP method, which has limited related work, or extracted through additional measurements to obtain the necessary channel model parameters.
@@
Figure 2 depicts the proposed simulation process extended by the 3GPP channel model[13, 14, 15]. In Step 1, users are required to perform general configurations, including antenna configuration, scenario settings, target types, and the positions of the BS, UT, and targets. Steps 2 to 5 are dedicated to large-scale parameter generation, while Steps 6 to 12 concentrate on small-scale parameter generation. Finally, Steps 13 and 14 are responsible for generating the channel coefficients. Steps 4, 9, and 12 correspond to large-scale parameter cascading, RCS generation, and small-scale parameter cascading, respectively. Step 15 refers to the combination of target and background channels.
2.1 Concatenate channel modeling
@@
In this section, We propose a parameter cascading method for ISAC channel simulation with the scenario layouts, path loss formulas, line-of-sight (LOS) probabilities, and large-scale parameters defined in the 3GPP channel model standards TR38.901, TR36.777, and TR37.885. Specifically, TR38.901 is employed for scenarios involving human and AGV sensing targets, TR36.777 is designated for UAV target scenarios, and TR37.885 is applied to V2X target scenarios. The cascading simulation method comprises two main components: large-scale parameter cascading and small-scale parameter cascading.
2.1.1 Large scale parameters concatenate
@@
As illustrated in Figure 2, the simulation commences with the establishment of various parameters, including the simulation scenario, target type, target position, movement speed and direction of the target, the number and location of base stations and terminals, antenna patterns, orientation angles of both base stations and terminals, system center frequency, and movement speed and direction of the terminals. For the configured scenario, the Line-of-Sight (LOS) probability and path loss for the Tx-Target and Target-Rx links are computed individually. The corresponding mean value of RCS is generated for various targets. The concatenation path loss can be calculated as
@@
where, and represent the distances of the Tx-target and target-Rx, respectively. denotes the center frequency of the simulation. is the speed of light with m/s. is the mean RCS.
2.1.2 Small scale parameters concatenate
The total expression for channel parameters generation utilized by the BUPTCMCC-6G-CMG+ channel simulator is as
Here, represents the power of the -th cluster, generated in Step 7; denotes the number of sub-paths within each cluster; and are the field components of the receiving antenna in the and directions, respectively, while and are the field components of the transmitting antenna in the and directions, respectively. P is the polarization matrix. and are the vertical and horizontal angles of the -th sub-path within the -th cluster, generated in Step 8. and are the unit vectors in spherical coordinates, with the calculation expressions as follows
where, and represent the position vectors of the receiving and transmitting antenna elements, respectively. is the linear value of the cross-polarization power ratio, generated in Step 10. is the wavelength corresponding to the carrier frequency band. and the Doppler frequency factor is calculated as
where, and represent the velocity vectors of the target and the receiver, respectively, and is the unit vector in spherical coordinates at the target’s position can be expressed as
Considering that the cascaded link between the Tx-target and target-Rx involves four possible scenarios: , , , and denoted as radio propagation case 1, 2, 3, and 4, respectively. Equation (2.1.2) presents the general channel expression under the NLOS-NLOS condition. When LOS conditions are involved, the corresponding LOS components in equation 2.1.2, such as power , angles and , and cross-polarization power ratio, should be selectively modified according to the expressions for the LOS case in the GBSM model. Let and be the number of clusters in the Tx-target and target-Rx links, respectively. The total number of clusters in the target channel is . For the n-th cluster of the target channel where ()
Cluster delay:
where and represent the delay of the p-th and q-th cluster from Tx-target and target-Rx links respectively.
Cluster power:
where and represent the power of the p-th and q-th cluster respectively.
Doppler frequency:
where and represent the doppler frequency of p-th and q-th links respectively.
Departure angle:
Arrival angle:
Polarization matrix P:
@@ 2.1.2 Small scale parameters concatenate
where and are the cross-polarization power ratio (XPR) in the Tx-target and target-Rx link respectively. S represents the polarization matrix of the target RCS introduced in section 2.2.1.
@@ 2.2 RCS modeling
where represents the distance resolution cell, is the speed of light, and is the pulse width. is the antenna beam solid angle, and is the distance from the radar receiver to the target. If the target’s size is contained within the volume , the target is considered a point target. For the five types of sensing targets defined by 3GPP UAVs, humans, long-distance vehicles, AGVs, and small road obstacles modeling as point targets is appropriate. However, for targets that exceed the resolution cell and possess irregular shapes, such as close-range vehicles and AGVs, large obstacles, and humans in gesture recognition scenarios, multi-point RCS modeling may be considered.

### v4

@@ Abstract
Integrated sensing and communication (ISAC) has been recognized as the key technology in the vision of the sixth generation (6G) era. With the emergence of new concepts in mobile communications, the channel model is the prerequisite for system design and performance evaluation. Currently, 3GPP Release 19 is advancing the standardization of ISAC channel models. Nevertheless, a unified modeling framework has yet to be established. This paper provides a simulation diagram of ISAC channel modeling extended based on the Geometry-Based Stochastic Model (GBSM), compatible with existing 5G channel models and the latest progress in the 3rd Generation Partnership Project (3GPP) standardization. We first introduce the progress of the ISAC channel model standardization in general. Then, a concatenated channel modeling approach is presented considering the team’s standardization proposals, which is implemented on the BUPTCMCC-6G-CMG+ channel model generator. We validated the model in cumulative probability density function (CDF) in statistical extension of angle and delay, and radar cross section (RCS). Simulation results show that the proposed model can realistically characterize the feature of channel concatenation and RCS within the ISAC channel.
@@
Channel model generator, channel concatenation, ISAC channel, RCS, 3GPP standardization.1 Introduction
@@
The sixth generation (6G) mobile communication technology is anticipated to enhance 5G communication capabilities roundly, including high capacity, low latency, and high reliability, while achieving native intelligence, integrated sensing and communication (ISAC), and ubiquitous connectivity [1]. The boundary between the physical and digital realms is expected to become increasingly indistinct, offering users a ubiquitous and immersive experience [2]. Concerns of communication systems will need to expand from information transmission to information acquisition and information processing[3]. ISAC is the essential technology necessary to achieve the information paradigm transformation described above. ISAC technology can facilitate the implementation of diverse use cases, particularly in localization, tracking, imaging, and the detection of target type, velocity, and direction. These capabilities will enable applications in autonomous driving, industrial automation, digital twin reconstruction of physical environments, gesture detection, and activity recognition[4, 5, 6, 7, 8]. The sensing information obtained from ISAC can facilitate the development of smart homes, factories, smart cities, and intelligent transportation systems. Meanwhile, ISAC technology introduces new performance metrics relative to traditional communication technologies, including detection probability, sensing capacity density, sensing accuracy, and resolution for velocity, angle, and delay [9].
@@
The transmission medium between the transmitter and receiver establishes the performance limits of communication systems. The wireless channel is crucial for the research and evaluation of wireless communication technologies [10, 11, 12]. The performance evaluation of the ISAC system introduces new requirements for channel models, particularly regarding sensing modes and scenarios, sensing use cases, and channel modeling methodologies. Considering sensing modes, current channel models usually utilize Bi-static modes for communication. Sensing applications necessitate considering the case of Mono-static channel models for self-sensing. Thus, ISAC channel models should incorporate both Mono-static and Bi-static sensing modes. Secondly, concerning sensing use cases involving sensing targets, ISAC applications expand the evaluation criteria to include target detection, localization, identification, and tracking. Examples encompass intrusion detection, unmanned aerial vehicle (UAV) tracking, vehicle-to-everything (V2X) communications, and the identification of diverse target types and sizes. That requires utilizing the scattering properties of radar cross section (RCS), polarization characteristics, delay, and Doppler information of electromagnetic waves as they interact with the target. Consequently, channel models should integrate the RCS characteristics of various targets. Moreover, for modeling methodologies, it is imperative to account for the concatenated characteristics of channels interacting with the target while coupling with existing standard communication channel models [13, 14, 15, 16].
@@
So far, the standardization of ISAC channel models by the 3rd Generation Partnership Project (3GPP) has held five conferences from RAN1 #116 to RAN1 #118-bis, wherein ISAC scenarios, types of sensing targets, sensing modes, and general channel model frameworks have been defined. Recent studies have demonstrated the latest progress for 3GPP ISAC channel modeling standardization [17]. The defined scenarios encompass traditional urban micro (UMi), urban macro (UMa), rural macro (RMa), and indoor hotspot (InH) scenarios, in addition to V2X, UAV, and industrial Internet of things (IIoT) scenarios. Accordingly, sensing targets are classified into five categories: humans, vehicles, UAVs, automated guided vehicles (AGVs), and obstacles, as shown in Table 1. Sensing modes are categorized into Bi-static and Mono-static, which are further subdivided into six specific use cases based on whether the sensing device is a user terminal (UT) or a base station (BS): BS-UT-Bi, BS-BS-Bi, UT-UT-Bi, UT-BS-Bi, UT-UT-Mono, and BS-BS-Mono [18]. In the framework of channel modeling, the ISAC channel is defined as the combination of the target channel and the background channel. The target channel includes all channel components that interact with targets, whereas the background channel comprises all other elements that do not engage with targets.
@@
For ISAC channel simulation, a channel model that separately generates communication and sensing clusters based on the 3GPP model and utilizes sensing echoes for sensing has been proposed [19]. Some studies introduce the concept of shared clusters to explain the phenomenon of scattering body reuse between the communication and sensing channels [5, 20, 21]. Some studies focus on deterministic ISAC channel modeling [22, 23]. However, these studies do not align with the target and background channel modeling framework or the concatenated channel method in 3GPP Rel-19. Most of these approaches use Mono-static echoes for sensing and do not support Bi-static sensing modes.
@@
To the authors’ knowledge, no unified ISAC channel simulation framework compatible with the latest 3GPP standardization progress has been proposed. Considering that standardization is primarily driven by industry and the existing standardization framework has not yet been fully finalized, proposing a unified channel expression that integrates key aspects discussed in the RAN1 agenda such as RCS modeling, the concatenated channel methodology, and micro-Doppler and macro-Doppler characteristics prior to 3GPP’s finalization is crucial for understanding the progress in ISAC standardization and the practical implementation of ISAC channel models. Given this, we present our solution, which is grounded in our team’s contributions to ISAC standardization and our understanding of standardized channel simulation. The contributions and innovations of this paper are outlined as follows:
Different from existing ISAC channel simulation models [19, 20, 21, 22, 23], we propose a unified ISAC channel simulation framework based on the 3GPP Rel-19 ISAC channel fundamental framework, where the ISAC channel is the combination of the target and background channels. This framework maximizes the use and compatibility with existing 5G standardized channel models. This work integrates our team’s contributions to the standardization process, presenting a solution that aligns with the latest standardized progress and consensus in ISAC channel modeling [17].
The effective ISAC channel simulation method has been validated.
We compared and analyzed the down-selection simulation results for the 8 concatenation options presented in the RAN1 #118-bis meeting, explaining the mechanisms behind the results. Based on it, we propose a simulation method for ISAC concatenated channels.
A unified RCS simulation model is introduced, which considers the progress of the RAN1 #116-118bis meetings and incorporates our team’s RCS measurement results. This model applies to human, vehicle, and UAV targets, accounting for factors such as large-scale, small-scale, statistical, and deterministic components.
The paper is organized as follows. Section describes the simulation architecture in detail. Section reports on evaluation metrics for ISAC. Then, Section presents the simulation results for both ISAC channel concatenation and RCS. Section concludes the paper.
2 Simulation Architecture
Building upon the current advancements in 3GPP standardization efforts and the GBSM simulation framework, this paper proposes an ISAC channel simulation framework designed for future standardization, which is integrated into the BUPTCMCC-6G-CMG+ channel model generator. The key characteristics of ISAC channels—including concatenated features, RCS polarization, RCS scattering properties, and the combination of target and background channels, are considered to establish a unified model for ISAC channel simulation. In this model, the ISAC channel is divided into two components: the target channel and the background channel. The geometric schematic of the model is shown in Figure 1.
@@ 2 Simulation Architecture
The target channel is further classified into Tx-target and target-Rx sub-channels. For Bi-static, the target channel is constructed by cascading the Tx-target and target-Rx sub-channels, with parameters for each sub-channel generated according to the existing 3GPP models. For Mono-static, the Tx-target link is generated using the 3GPP model, whereas the target-Rx components are derived by utilizing channel reciprocity. Furthermore, the background channel is generated directly using the 3GPP statistical channel model for the Bi-static sensing mode. As to the Mono-static sensing mode, the background channel can be derived from the 3GPP method, which has limited related work, or extracted through additional measurements to obtain the necessary channel model parameters.
@@
Figure 2 depicts the proposed simulation process extended by the 3GPP channel model[13, 14, 15]. In Step 1, users can perform general configurations, including antenna configuration, scenario settings, target types, and BS, UT, and target positions. Steps 2 to 5 are dedicated to large-scale parameter generation, while Steps 6 to 12 concentrate on small-scale parameter generation. Finally, Steps 13 and 14 are responsible for generating the channel coefficients. Steps 4, 9, and 12 correspond to large-scale parameter concatenation, RCS generation, and small-scale parameter concatenation. Step 15 refers to the combination of target and background channels.
2.1 Concatenated channel modeling
@@
In this section, we propose a parameter-concatenated method for ISAC channel simulation with scenario layouts, path loss formulas, line-of-sight (LOS) probabilities, and large-scale parameters defined in the 3GPP channel model standards TR38.901, TR36.777, and TR37.885. Specifically, TR38.901 is used for scenarios involving human and AGV sensing targets, TR36.777 is designated for UAV scenarios, and TR37.885 is applied to V2X scenarios. The concatenated simulation method comprises two main components: large-scale parameters concatenation and small-scale parameters concatenation.
2.1.1 Large scale parameters concatenation
@@
As illustrated in Figure 2, the simulation starts with the establishment of various parameters, including the simulation scenario, target type, target position, velocity and direction, the number and location of base stations and terminals, antenna patterns, orientation of both base stations and terminals, system center frequency, speed and direction of the terminals. For the configured scenario, the LOS probability and path loss for the Tx-target and target-Rx links are computed individually. The corresponding mean value of RCS is generated for the target. The concatenation path loss can be calculated as
@@
where and represent the distances of the Tx-target and target-Rx, respectively. denotes the center frequency of the simulation. is the speed of light with m/s. is the mean RCS.
2.1.2 Small scale parameters concatenation
In this section, we derive and present a unified function for the ISAC target channel. This function is compatible with the concatenated channel methodology outlined in the 3GPP RAN1 ISAC channel modeling issue, incorporating key parameters such as RCS small-scale variations, RCS polarization, and Doppler. The unified function for each receiver and transmitter antenna element pair u and s in LOS+LOS condition is given in (2.1.2), while the equations for the remaining LOS+NLOS, NLOS+LOS, and NLOS+NLOS conditions can be obtained by simplifying 2.1.2.
Here, and represent the Ricean K-factors for the Tx-target and target-Rx links, respectively, which are generated in Step 5 and converted to a linear scale. denotes the Dirac delta function. is the sum of the delays of LOS rays in Tx-target and target-Rx links. The terms , , , , and correspond to the matching components for LOS-LOS path, LOS-NLOS multipath, NLOS-LOS multipath, and NLOS-NLOS multipath of the target channel coefficients respectively. Their detailed expressions are provided in (2.1.2), (2.1.2), (2.1.2), and (6).
Here, and are the number of clusters in the Tx-target and target-Rx links, respectively. is the wavelength corresponding to the center carrier frequency. and are the number of rays in each cluster in the Tx-target and target-Rx links, respectively. and represent the field components of the receiving antenna in the zenith and azimuth directions, respectively. and represent the zenith and the azimuth angle, respectively. and represent the field components of the transmitting antenna in the zenith and azimuth directions, respectively. represents the RCS value related to angular terms, the detailed definition of which can be found in Section 2.2.2. and are the power of the -th cluster in the Tx-target link and the power of the -th cluster in the target-Rx link, respectively. and represent the position vectors of the receiving and transmitting antenna elements, respectively. and are the unit vectors in spherical coordinate, with the calculation expressions as follows:
is the Doppler frequency for a target including both macro-Doppler and micro-Doppler, which can be modeled using a unified formula as:
where, is the spherical unit vector at the receiver for the link from Rx to the scattering point. is the spherical unit vector at the scattering point for the link from the scattering point to Rx. is the spherical unit vector at the receiver for the link from Tx to the scattering point. is the spherical unit vector at the scattering point for the link from the scattering point to Tx. , , and are the velocity of the Tx, Rx, and scattering point respectively. can include macro-Doppler and micro-Doppler motion, .
, , , and are polarization matrices correspond to the LOS-LOS, LOS-NLOS, NLOS-LOS, and NLOS-NLOS path of the target channel, respectively.
@@ 2.1.2 Small scale parameters concatenation
Where, and are the cross-polarization power ratio (XPR) in the Tx-target and target-Rx link, respectively. and represent the phase of the LOS ray of Tx-target and target-Rx links, respectively. Here, and are the three dimension (3D) distance of Tx-target and target-Rx links, respectively. S is the polarization matrix of the target RCS introduced in section 2.2.1.
@@ 2.2 RCS modeling
where represents the distance resolution cell, and is the pulse width. is the antenna beam solid angle, and is the distance from the radar receiver to the target. If the target’s size is contained within the volume , the target is considered a point target. For the five types of sensing targets defined by 3GPP UAVs, humans, long-distance vehicles, AGVs, and small road obstacles modeling as point targets is appropriate. However, for targets that exceed the resolution cell and possess irregular shapes, such as close-range vehicles and AGVs, large obstacles, and humans in gesture recognition scenarios, multi-point RCS modeling may be considered.